# Learning Logical Rules using Minimum Message Length

### v1

@@ Introduction
This learner searches for programs which minimise a user-specified cost function. In this paper, this cost is either C-MDL or an approximation of MML. The ApproxLearner learning procedure is similar to the inductive ILP system MaxSynth, which minimises the C-MDL cost function (Hocquette et al. 2024). Both procedures (i) generate individual rules, (ii) test and cache the size and performance statistics of each rule on training data, and (iii) find good combinations of these rules (programs) according to the cost. We describe these stages next.
@@
Constraints: The main difference between ApproxLearner and MaxSynth is step (i). When generating rules, MaxSynth uses C-MDL-specific constraints to efficiently prune rules from the rule space that cannot be in a C-MDL optimal hypothesis. This pruning vastly shrinks the search space and enables MaxSynth to learn globally optimal programs efficiently. Conversely, ApproxLearner uses weaker, cost function independent constraints (pruning generalisations of inconsistent rules) and is much less efficient. In the absence of MML-specific constraints, we use the ApproxLearner procedure to compare MML and C-MDL costs fairly, leaving MML-specific constraints to future work.
@@
Search: To implement the search for good rule combinations in step (iii), both MaxSynth and ApproxLearner use constraint solvers, specialised tools which efficiently search large combinatorial spaces given a cost to optimise. ApproxLearner uses the CP-SAT (Perron and Didier 2025) constraint solver. CP-SAT requires a linear cost to perform an effective search. The C-MDL cost is already linear and is modelled directly in CP-SAT. The MML cost is not linear, so we use a linear approximation.
@@
Linear MML Approximation: We cache the values of the length of the rule-specific information $C_{R^{\prime}}(H)$ (equation 7) for each rule, as it behaves linearly: $C_{R^{\prime}}(H_{1}\cup H_{2})=C_{R^{\prime}}(H_{1})+C_{R^{\prime}}(H_{2})$ for any two hypotheses $H_{1},H_{2}$. For all other terms in the $C(H,E)$ sum (equation 2), we select five breakpoints and linearly interpolate between each to determine a piecewise linear (PL) approximation. We provide the PL approximations, the cached $C(R)$ values, and rule statistics to the CP-SAT solver so that it may implement the search procedure.
Herbrand Base Calculations
@@
For efficiency, in the RandomLearner experiments, the Herbrand Base is calculated once from the original trial and reused for any trial variation (altering training example size, noise level, or example balance). In the ApproxLearner experiments (Q6 below), however, the Herbrand Base is recalculated for all trials. The Herbrand Base is used to calculate the generality predicate priors and the set of possible example literals, which are all literals in the Herbrand Base that contain the example relation.
Experiments
@@
We answer Q1 by comparing the testing accuracy of MML and C-MDL programs on 30 diverse datasets.
@@
We answer Q2 by comparing the testing accuracy of MML and C-MDL programs on datasets with different proportions of examples, ranging from 10:0 (100% positive examples), 8:2, 6:4, …, 0:10 (100% negative examples). We randomly generate these datasets from those used in Q1. We restrict dataset sizes to 20 and 50 to control for the effect of the number of training examples.
@@
We answer Q4 by comparing the testing accuracy of our standard MML procedure (with generality predicate priors) against MML with uniform predicate priors on the same datasets as Q3.
@@
    Can MML handle similar amounts of noise as C-MDL?
We answer Q5 by introducing different amounts (10%, 20%, … 50%) of random noise into each of the 30 datasets and comparing the testing accuracies of the MML and C-MDL selected programs. We also run MML vs C-MDL on the real-world noisy alzheimer datasets. We introduce noise by randomly selecting a subset of training examples, which we randomly reassign as positive or negative.
  1. Q6
@@
We answer Q6 by comparing C-MDL and MML on the noiseless datasets in Q1 and on modifications of these datasets containing only the positive or negative examples using ApproxLearner.
Experimental Choices
@@
Theta Priors We use identical beta prior distributions with $\alpha$ and $\beta$ parameters for our $\theta_{+}$ and $\theta_{-}$ values chosen to ensure that the learned $\theta$ values are always greater than $0.5$. This constraint has the following effect. When $\theta_{+}>0.5$, if our learned program holds, the relation that we describe is also likely to hold. When $\theta_{-}>0.5$, if our learned program does not hold, the relation that we describe is also likely not to hold. Constraining $\theta_{+},\theta_{-}>0.5$ results in the standard, non-negated program interpretation. The expected $\theta$ value is $e=\frac{\alpha}{\alpha+\beta}$. The probability of error $r$, used in $C(H_{+/-})$ calculation (Equation 3) is r=1er=1-e. For noiseless tasks, we assume $r$ is near zero: $\alpha=1000,000$ and $\beta=1$. For noisy tasks, we assume weaker, but still discriminatory programs: $\alpha=5000$, $\beta=1$ . When $\beta=1$, choosing $\alpha>N+1$, where $N$ is the number of training examples, enforces our constraint, which is satisfied for our largest considered training dataset (<2000<2000 examples).
@@
Alzheimer. Real-world noisy datasets, learning attributes of drug molecules for alzheimers treatment (King et al. 1992).
Zendo. An inductive game where the goal is to determine an underlying true rule (Cropper and Hocquette 2023).
1D-ARC. A one-dimensional adaptation (Xu et al. 2023) of the ARC-AGI dataset (Chollet 2019).
Trains. Synthetic datasets whose goal is to determine a rule which differentiates eastbound from westbound trains (Larson and Michalski 1977).
Results
@@
Q1: Overall Performance: RandomLearner
@@
Q2: Example Balance: RandomLearner
For training sets of size 20 and 50, the mean pairwise balanced accuracy difference consistently favours MML across all 12 tested size-example proportion datasets. This advantage is statistically significant ($p<0.05$) in 11/12 conditions. Figure 1 illustrates this trend for the size 20 datasets. One reason why MML outperforms C-MDL is because C-MDL struggles when learning from 100% positive examples. In this setup, a program that classifies all examples as true is C-MDL optimal (it makes no errors and is minimal in size). By contrast, the MML likelihood (Equation 1) penalises overly general programs. Therefore, a program which entails all possible examples results in the cost of Equation Literals as $\log\binom{|T|}{tp+fp}$ which, when $|T|>>tp+fp$, is very large. MML prefers a program which entails only the seen examples (and makes no/few errors), as this cost is instead $\log\binom{tp+fp}{tp+fp}=0$. These results support Claim 2.
Q3: Data Efficiency: RandomLearner
@@
Q4: Generality Priors: RandomLearner
@@
Q5: Noise Tolerance: RandomLearner
On datasets with 10-50% added noise or the standard alzheimers datasets, we find no statistically significant difference between MML and C-MDL. We hypothesise that MML’s performance in these low-signal domains is hindered by a prior that is overly optimistic, and would be improved by one which is more carefully selected. These results support Claim 5.
Q6: Example Balance Tolerance and Overall Performance: ApproxLearner
On datasets consisting of exclusively positive and exclusively negative examples, the approximate MML cost significantly ($p<0.01$) outperforms C-MDL by $9.2\pm 2.7$% and $25.9\pm 4.0\%$ respectively. The maximum improvement was $50\%$ in both experiments: the maximum possible. The maximum losses were much smaller, $3.6\%$ and $0.0\%$ for $100\%$ and $0\%$ positive examples, respectively. Experiments on the standard noiseless tasks reveal no significant difference between MML and C-MDL. This result is despite the difficulty of optimising the more complex MML cost function. These results support Claim 6
Conclusions and Limitations
We derived an ILP cost function that is data-efficient, insensitive to example imbalance, and noise-tolerant. Our cost function enables learning from exclusively positive examples, a common failure mode for many ILP systems. Empirically, our cost outperforms the score from the state-of-the-art MDL system MaxSynth on data efficiency and example imbalance, and matches its performance on noisy data. Our introduced prior preference for more general predicates results in significantly increased data efficiency relative to a uniform prior when learning from a single example. Unlike MaxSynth, our system does not guarantee optimal programs. Future work should identify tight constraints to enable efficient optimal search to enable equivalently powerful MML-based systems. The described approach has only been demonstrated for finite domains, but may be extended to infinite domains by constructing a truncated Herbrand Base from only the ground terms observed. Alternative priors should also be explored, such as extending the presented predicate generality priors to incorporate domain knowledge. We acknowledge that the beta prior also requires user specification, which is not needed for MDL. One should instead determine robust noisy default beta priors by grid search over many datasets. This should improve performance, particularly in high-noise domains.
Acknowledgements
Ruben Sharma was supported by the EPSRC Centre of Doctoral Training in Automated Chemical Synthesis Enabled by Digital Molecular Technologies (Syntech, EP/S005722/1). Sebastijan Dumančić was supported by NWO (OCENW.M.21.325 and NWA.1418.22.008). Ross D. King was supported by the EPSRC (EP/R022925/2, EP/W004801/1 and EP/X032418/1), and by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Alice Wallenberg Foundation. Andrew Cropper was supported by his EPSRC fellowship (EP/V040340/1).

### v2

@@ 6.2 ApproxLearner
This learner searches for hypotheses which minimise a user-specified cost function. In this paper, this cost is either C-MDL or an approximation of MML. The ApproxLearner learning procedure is similar to the inductive ILP system MaxSynth, which minimises the C-MDL cost function Hocquette et al. (2024). Both procedures (i) generate individual rules, (ii) test and cache the size and coverage of each rule on training data, and (iii) find good combinations of these rules according to the cost. As in RandomLearner, we assess a deterministic rule or combination of rules by treating it as $\text{det}(h)$ and converting to $h$ (definition 2) for scoring. Pseudocode for this procedure is shown in Algorithm 1. We describe these stages next.
def ApproxLearner(bk, pos, neg, max_size):
bst_sol, bst_score, size = {}, $\infty$$, 1
poss = make_poss_exs(bk,pos,neg)
seen_hyps, cons = {}, {}
while size $\leq$$ max_size:
h = generate(cons, size)
if h == UNSAT:
size += 1
continue
tp,fn,tn,fp,gen = test(pos,neg,bk,poss,h)
h_score = score(h,tp,fp,tn,fn,gen)
if h_score < bst_score:
bst_sol, bst_score = h, h_score
seen_hyps += h
comb = combine(seen_hyps, bst_score)
if comb != UNSAT:
bst_sol = comb
test_res = test(pos,neg,bk,poss,comb)
tp,fn,tn,fp,gen = test_res
bst_score=score(comb,tp,fp,tn,fn,gen)
cons += constrain(h,fp,fn)
return bst_sol
Algorithm 1 ApproxLearner algorithm
@@
Constraints: The main difference between ApproxLearner and MaxSynth is step (i). When generating rules, MaxSynth uses C-MDL-specific constraints to efficiently prune rules from the rule space that cannot be in a C-MDL optimal hypothesis. This pruning vastly shrinks the search space and enables MaxSynth to learn globally optimal hypotheses efficiently. Conversely, ApproxLearner uses weaker, cost function independent constraints (such as pruning generalisations of inconsistent rules) and is much less efficient. In the absence of MML-specific constraints, we use the ApproxLearner procedure to compare MML and C-MDL costs fairly, leaving MML-specific constraints to future work.
@@
Search: To implement the search for good rule combinations in step (iii), both MaxSynth and ApproxLearner use constraint solvers, specialised tools which efficiently search large combinatorial spaces given a cost to optimise. ApproxLearner uses the CP-SAT Perron and Didier constraint solver. CP-SAT requires a linear cost to perform an effective search. The C-MDL cost is already linear and is modelled directly in CP-SAT. The MML cost is not linear, so we use a linear approximation.
@@
Linear MML Approximation: We cache the values of all message costs that behave linearly. For all other terms that compose the $C(h,E)$ sum (equation 1), we select five breakpoints and linearly interpolate between each to determine a piecewise linear (PL) approximation. We provide the PL approximations, the cached rule encoding values, and rule statistics to the CP-SAT solver so that it may implement the search procedure.
6.2.1 Herbrand Base Calculations
@@
For efficiency, in the RandomLearner experiments, the Herbrand Base is calculated once from the original trial and reused for any trial variation (altering training example size, noise level, or example balance). In the ApproxLearner experiments (Q5 below), however, the Herbrand Base is recalculated for all trials.
7 Experiments
@@ 1. Q1a
We answer Q1a by comparing the testing accuracy of MML and C-MDL programs on 30 diverse datasets.
  1. Q1b
    Can MML handle similar amounts of noise as C-MDL?
We answer Q1b by introducing different amounts (10%, 20%, … 50%) of random noise into each of the 30 datasets and comparing the testing accuracies of the MML and C-MDL selected programs. We also run MML vs C-MDL on the real-world noisy alzheimer datasets. We introduce noise by randomly selecting a subset of training examples, which we randomly reassign as positive or negative.
@@ 1. Q1b
We answer Q2 by comparing the testing accuracy of MML and C-MDL programs on datasets with different proportions of examples, ranging from 10:0 (100% positive examples), 8:2, 6:4, …, 0:10 (100% negative examples). We randomly generate these datasets from those used in Q1a. We restrict dataset sizes to 20 and 50 to control for the effect of the number of training examples.
@@
We answer Q4 by comparing the testing accuracy of our standard MML procedure (with generality predicate priors) against MML with uniform predicate priors on the same datasets as Q2.
@@
We answer Q5 by comparing C-MDL and MML on the noiseless datasets in Q1 and on modifications of these datasets containing only the positive or negative examples using ApproxLearner.
7.1 Experimental Choices
@@
Theta Priors The expected $\theta_{+}$ and $\theta_{-}$ values are $e=\frac{\alpha}{\alpha+\beta}$. The probability of misclassification error $r$, used in $C_{cov}(H)$ calculation is r=1er=1-e. For noiseless tasks, we assume $r$ is near zero: $\alpha=1000,000$, $\beta=1$. For noisy tasks, we assume weaker, but still discriminatory programs: $\alpha=5000$, $\beta=1$. The beta priors chosen ensure $\theta_{+},\theta_{-}>0.5$.
@@ 7.1.1 Datasets
Alzheimer. Real-world noisy datasets, learning attributes of drug molecules for alzheimers treatment King et al. (1992).
Zendo. An inductive game where the goal is to determine an underlying true rule Cropper and Hocquette (2023).
1D-ARC. A one-dimensional adaptation Xu et al. (2023) of the ARC-AGI dataset Chollet (2019).
Trains. Synthetic datasets whose goal is to determine a rule which differentiates eastbound from westbound trains Larson and Michalski (1977).
7.2 Results
@@
7.2.1 Q1a: Overall Performance: RandomLearner
@@
7.2.2 Q1b: Noise Tolerance: RandomLearner
On datasets with 10-50% added noise or the standard alzheimers datasets, we find no statistically significant difference between MML and C-MDL. We hypothesise that MML’s performance in these low-signal domains is due to a prior that is overly optimistic, and would be improved by one which is more carefully selected. These results support Claim 1.
7.2.3 Q2: Example Balance: RandomLearner
For training sets of size 20 and 50, the mean pairwise balanced accuracy difference consistently favours MML across all 12 tested size-example proportion datasets. The advantage is statistically significant ($p<0.05$) in 11/12 conditions. Figure 1 illustrates this trend for the size 20 datasets. One reason why MML outperforms C-MDL is because C-MDL struggles when learning from 100% positive examples. In this setup, a program that classifies all examples as true is C-MDL optimal (it makes no errors and is minimal in size). By contrast, the MML likelihood (Equation 4) penalises overly general programs. A program that entails all possible examples results in the cost of $\text{atoms}(E|h)$ as $\log\binom{|T|}{tp+fp}$ which, when $|T|>>tp+fp$, is very large. MML prefers a program which entails only the seen examples (and makes no/few errors), as this cost is instead $\log\binom{tp+fp}{tp+fp}=0$. These results support Claim 2.
7.2.4 Q3: Data Efficiency: RandomLearner
@@
7.2.5 Q4: Generality Priors: RandomLearner
@@
7.2.6 Q5: Example Balance Tolerance and Overall Performance: ApproxLearner
On datasets consisting of exclusively positive and exclusively negative examples, the approximate MML cost significantly ($p<0.01$) outperforms C-MDL by $9.2\pm 2.7$% and $25.9\pm 4.0\%$ respectively. The maximum improvement was 50% in both experiments: the maximum possible. The maximum losses were much smaller, $4.8\%$ and $0.0\%$ for $100\%$ and $0\%$ positive examples, respectively. Experiments on the standard noiseless tasks reveal no significant difference between MML and C-MDL. This result is despite the difficulty of optimising the more complex MML cost function. These results support Claim 5
8 Conclusions and Limitations
We derived an ILP cost function that is data-efficient, insensitive to example imbalance, and noise-tolerant. Our cost function enables learning from exclusively positive examples, a common limitation for many ILP systems Cropper and Morel (2021); Hocquette et al. (2024). Empirically, our cost function outperforms the that of the state-of-the-art MDL system MaxSynth on data efficiency and example imbalance, and matches its performance on noisy data. Our prior preference for more general predicates results in significantly increased data efficiency relative to a uniform prior when learning from a single example.
Limitations
Unlike MaxSynth, our system does not guarantee optimal programs. Future work should identify better constraints for ApproxLearner to enable efficient optimal search for MML-based systems. The described approach has only been demonstrated for finite domains, but may be extended to infinite domains by constructing a truncated Herbrand Base from only the ground terms observed. Alternative priors should also be explored, such as extending the presented predicate generality priors to incorporate domain knowledge. The beta prior requires user specification, which is not needed for MDL. One should instead determine robust noisy default beta priors by grid search over many datasets, which should improve performance, particularly in high-noise domains.